OpenStudy (anonymous):

In a sports equipment locker, there are 81 balls and 45 baseball mitts. Write the ratio of balls to baseball mitts in simplest form

OpenStudy (anonymous):

81/45. Simplify that. 9/5

OpenStudy (anonymous):

\[\frac{ 81 }{ 45 } = \frac{ 9 \times 9 }{ 5 \times 9} = \frac{ 9 }{ 5 } \times \frac{ 9 }{ 9 } = \frac{ 9 }{ 5 }\]